OpenStudy (anonymous):

Jonathan collects postcards and stamps. The number of postcards in his collection is 12 more than 3/4 the number of stamps. He has 39 postcards in all. If Jonathan has x stamps, which equation represents this situation and how many stamps does he have? A) The equation is 3/4x +12=39 , and the number of stamps is 36. B) The equation is 3/4x +x=39 , and the number of stamps is 24. C) The equation is x+3/4+x=39 , and the number of stamps is 18. D) The equation is 3/4x= 39+x , and the number of stamps is 156.

OpenStudy (anonymous):

just simply translate the sentence into equation form 12 + (3/4)x = 39 Put x to a side by itself (3/4)x = 39 - 12 x = 27/ (3/4) = 27 x (4/3) x = 36 Answer = A